OpenStudy (anonymous):

5+1/2 multiply 5+1/2

OpenStudy (lgbasallote):

do you mean \[\huge 5 + \frac 12 \times 5 + \frac 12\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

hint: according to PEMDAS multiply 1/2 x 5 first

OpenStudy (anonymous):

No. 5/1/2

OpenStudy (lgbasallote):

what is the original question?

OpenStudy (anonymous):



OpenStudy (lgbasallote):

you sure? the question wasn't originally \[\huge 5 + \frac 12 \times 5 + \frac 12?\] because those are two different things...

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

It is how I drew it.

OpenStudy (lgbasallote):

well then change both to improper fractions first \[\huge \implies \frac{11}2 \times \frac{11}2\] do you now what to do now?

OpenStudy (anonymous):

121/4?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

now change it to mixed fraction to see something cool

OpenStudy (anonymous):

UH O.O I got.. 32.5